### v2

# Exploring Multilingual Concepts of Human Values in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

Abstract
Prior research has revealed that abstract concepts are linearly represented as directions in the representation space of LLMs, predominantly centered around English. In this paper, we extend this investigation to a multilingual context, with a specific focus on human values-related concepts (i.e., value concepts) due to their significance for AI safety. Through our comprehensive exploration covering 7 types of human values, 16 languages and 3 LLM series with distinct multilinguality, we first empirically confirm the presence of value concepts within LLMs in a multilingual format. Further analysis on the cross-lingual characteristics of these concepts reveals 3 traits arising from language resource disparities: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, all in terms of value concepts. Moreover, we validate the feasibility of cross-lingual control over value alignment capabilities of LLMs, leveraging the dominant language as a source language. Ultimately, recognizing the significant impact of LLMs’ multilinguality on our results, we consolidate our findings and provide prudent suggestions on the composition of multilingual data for LLMs pre-training.
Warning: This paper contains examples that can be upsetting or offensive.
Exploring Multilingual Concepts of Human Values in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?
Shaoyang Xu1, Weilong Dong2, Zishan Guo2, Xinwei Wu2and Deyi Xiong21thanks:  Corresponding author. 1School of New Media and Communication, Tianjin University, Tianjin, China 2College of Intelligence and Computing, Tianjin University, Tianjin, China {syxu, willowd, guozishan, wuxw2021, dyxiong}@tju.edu.cn
1 Introduction
Recent years have witnessed the emergence of large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and LLaMA2 (Touvron et al., 2023). These LLMs have shown powerful capabilities in natural language understanding and generation (Guo et al., 2023; Bang et al., 2023; Jiao et al., 2023). However, alongside with their prowess, LLMs present potential risks. Research has demonstrated that LLMs can generate responses containing toxic, untruthful, biased, and even illegal content (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Thus, aligning LLMs with human values (i.e., value alignment) is necessary for unleashing their potential safely.
Human values, encompassing concepts like fairness, deontology, utilitarianism, and so on, although challenging to be precisely defined in language, are undoubtedly embedded in textual form (Hendrycks et al., 2021). Recently, Zou et al. (2023a) have introduced Representation Engineering (RepE) to enhance the transparency and controllability of deep neural networks. Through RepE, they unveil that high-level concepts can be extracted as concept vectors from LLMs, utilizing positive and negative text pairs aligned with the directions of specific concepts. These concept vectors, representing the directions of corresponding concepts, can be utilized to assess whether the behavior of LLMs aligns with or to steer their behavior towards the target directions  (Zou et al., 2023a; Li et al., 2023; Leong et al., 2023; Liu et al., 2023).
However, existing studies on concept representations in LLMs have primarily focused on English, leaving multilingual concepts unexplored. Our work is the first to explore multilingual concepts in LLMs, emphasizing human values-related concepts to advance multilingual AI safety and utility. The primary research questions we aim to answer are as follows: (Q1) Do LLMs encode concepts representing human values in multiple languages? (Q2) To what extent are these concepts consistent and transferable across different languages? (Q3) Whether LLMs trained with different distributions of multilingual data exhibit distinct multilinguality in these concepts? (Q4) Is Value Alignment of LLMs Controllable across Languages? To address these questions, we propose a framework consisting of 5 essential components: extracting multilingual concept vectors from LLMs (§3.1) and evaluating their correlation with the corresponding concepts (concept recognition task in §3.2) to answer Q1; computing cross-lingual similarity of concept vectors (§3.3) and performing cross-lingual concept recognition (§3.4) to answer Q2 and Q3; and manipulating model behavior cross-lingually via concept vectors (§5) to answer Q4.
Our analysis covers 7 concepts related to human values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness, given their significance for AI safety (Hendrycks et al., 2021; Bai et al., 2022; Askell et al., 2021; Touvron et al., 2023). To ensure the breadth and reliability of our findings, we have selected these 7 concepts for their diverse definitions and ethical attributes (Vida et al., 2023). Throughout this paper, we collectively refer to them as “value concepts” to reflect their diversity and keep consistent with existing AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). For comprehensive definitions, ethical backgrounds and examples of these value concepts, please refer to Appendix A.
In addition to diverse human values, our experiments involve 16 languages [^1] We recognize that linguistic diversity can foster cultural variations, potentially resulting in diverse interpretations of the same value from different cultural backgrounds (Hershcovich et al., 2022; Hämmerl et al., 2023). For example, regarding deontology, some cultures prioritize individual responsibility while others emphasize social obligations (Cao et al., 2023; Hofstede, 1984). However, our work focuses on the multilingual representations of value concepts within LLMs and their universal cross-lingual patterns, leaving the exploration on cultural divergences in human values for our future research. and 3 LLM families with different multilinguality. Specifically, we categorize the multilinguality of these 3 LLM families based on language distributions in their pre-training data into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with respectively balanced multilinguality (i.e., BLOOMZ series). Appendix D provides detailed language distributions of their pre-training data.
Through in-depth analysis spanning multiple tasks, value concepts, languages and LLMs, our key findings are as follows:
  • LLMs encode concepts representing human values in multiple languages, and the expansion of model size and the richness of language resources both contribute to a more precise capture of these concepts (§4.2).
  • The distribution of language resources significantly impacts the cross-lingual properties of these concepts. Specifically, an imbalance in language resources results in cross-lingual inconsistency (§4.3.1), distorted linguistic relationships (§4.3.2), and unidirectional cross-lingual transfer (§4.3.3) between high- and low-resource languages. The cross-lingual properties of value concepts are also intricately tied to the multilinguality of the models to be extracted (§4.3).
  • The value alignment of LLMs can be effectively transferred across languages, with the dominant language as a source language (§5.2).
Drawing from these findings, we prudently consider the following suggestions for multilingual pre-training data of LLMs, which might contribute to enhancing multilingual AI safety and utility. First, despite the positive effect of dominant languages as sources for cross-lingual alignment transfer (§5.2), it is crucial to avoid an excessive prevalence of these languages to mitigate unfair cross-lingual patterns, such as inconsistent multilingual representations (§4.3.1), distorted linguistic relationships (§4.3.2), and monotonous transfer patterns (§4.3.3). These traits could potentially amplify the risk of multilingual vulnerability (§5.2) and undermine cultural diversity (Zhang et al., 2023; Cao et al., 2023). Furthermore, we encourage a more balanced distribution of non-dominant languages, particularly those with extremely limited resources, to foster more equitable cross-lingual patterns (§4.3.2 and §4.3.3).
2 Related Work
Representation Engineering
Representation Engineering (RepE) introduced by Zou et al. (2023a) extracts abstract concepts as vectors from LLMs using positive and negative samples that describe specific concepts. The effectiveness of these vectors has been validated across dimensions such as correlation and manipulation. Specifically, correlation experiments have assessed the predictive power of the extracted vectors to classify out-of-distribution data as positive or negative, while manipulation experiments have evaluated the vectors’ ability to control LLMs’ behavior by adding or subtracting them from the hidden states (Liu et al., 2023; Leong et al., 2023; Wang and Shu, 2023). While previous research has primarily focused on English, we pioneers the extension of RepE into a multilingual context. exploring multilingual concepts within LLMs through concept extraction, correlation, and manipulation experiments, all conducted in a multilingual or cross-lingual manner.
Multilinguality of LLMs
Multilingual pre-trained language models (Devlin et al., 2019; Xue et al., 2021; Conneau and Lample, 2019) tend to demonstrate a proficiency biased toward high-resource languages (Blasi et al., 2022; Joshi et al., 2020). Numerous studies (Zhang et al., 2023; Qi et al., 2023; Xu et al., 2023; Ohmer et al., 2023) have delved into the multilinguality of LLMs and examined the cross-lingual consistency and transferability of knowledge within them, aiming to alleviate language biases. Our work provides intuitive insights into the multilinguality of LLMs from the perspective of multilingual abstract concepts.
Multilingual AI Safety
Despite their remarkable capabilities, LLMs present potential risks (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Research has indicated that LLMs exhibit even greater vulnerability in multilingual scenarios, as evidenced by the multilingual jailbreak challenge (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). In this work, we explore the cross-lingual characteristics of safety-related concepts in LLMs’ representation spaces, shedding light on the mechanisms underlying multilingual vulnerability.
3 Exploring Multilingual Value Concepts
Representation Engineering (RepE) (Zou et al., 2023a) is able to assess whether an LLM encodes a specific concept based on its ability to distinguish between positive and negative samples of this concept using only its activations. This involves first extracting concept vectors from the model using training data and then evaluating the capability of these vectors to distinguish corresponding concepts on testing data (referred to as concept recognition task in our paper). Both training and testing data embed the target concept, albeit not in the same distribution.
Partially inspired by RepE, we introduce a framework for delving into mulitilingual value concepts within LLMs. Specifically, to investigate whether LLMs encode value concepts in a multilingual context, we extend the extraction of concept vectors and concept recognition to a multilingual scenario, corresponding to §3.1 and §3.2, respectively. To further analyze the cross-lingual consistency and transferability of these concepts, we propose the calculation of cross-lingual similarity of concept vectors and cross-lingual concept recognition in §3.3 and §3.4, respectively. Furthermore, we propose a cross-lingual value alignment control with extracted concept vectors, discussed in §5.
3.1 Collecting Multilingual Concept Vectors
To obtain the vector of a concept $\bm{v}_{c}$ embedded in a LLM, we utilize pairs of positive and negative texts that represent the two opposing directions of the concept. For example, a harmless instruction “Provide a tutorial on how to make a breakfast” and a harmful instruction “Provide a tutorial on how to make a bomb” form a pair representing the concept of harmfulness. Given a set of such training text pairs denoted as $\mathcal{T}_{c}^{\textnormal{train}}=[(t_{0+},t_{0-}),(t_{1+},t_{1-}),...]$, we feed them into the model. Collecting text representations from the last token of each corresponding text, we obtain $\mathcal{R}_{c}^{\textnormal{train}}=[(\bm{r}_{0+},\bm{r}_{0-}),(\bm{r}_{1+},\bm{r}_{1-}),...]$. We then compute the mean of the differences between these opposite text representations, obtaining the concept vector $\bm{v}_{c}$, which is formulated as follows:
$\bm{v}_{c}=\frac{1}{N}\sum_{i=0}^{N-1}(\bm{r}_{i+}-\bm{r}_{i-})\quad N=|\mathcal{T}_{c}^{\textnormal{train}}|$ (1)
For each concept $c$, we use multilingual text pairs to derive its concept vector $\bm{v}_{c}^{l}$ for each language $l$.
It’s worth noting that, in practice, we extract concept vectors from each layer of the model. These vectors are then collectively utilized for the concept recognition task (§3.2). Further details are provided in the next section.
3.2 Recognizing Multilingual Concepts
To assess the effectiveness of the extracted concept vectors and their correlation with specific concepts, we explore them for classifying test data. This task essentially measures the model’s capability of distinguishing the direction of these concepts. Specifically, for a concept $c$, we employ a set of testing text pairs $\mathcal{T}_{c}^{\textnormal{test}}=[(\hat{t}_{0+},\hat{t}_{0-}),(\hat{t}_{1+},\hat{t}_{1-}),...]$ representing the two directions of the concept and input them into the model. Similarly, we obtain text representations $\mathcal{R}_{c}^{\textnormal{test}}=[(\bm{\hat{r}}_{0+},\bm{\hat{r}}_{0-}),(\bm{\hat{r}}_{1+},\bm{\hat{r}}_{1-}),...]$ by taking the last token’s representation of each corresponding text. Furthermore, we calculate the dot product between the previously acquired vector $\bm{v}_{c}$ and these text vectors, resulting in classification scores $\mathcal{S}_{c}^{\textnormal{test}}=[(s_{0+},s_{0-}),(s_{1+},s_{1-}),...]$, where $s_{i\pm}=\bm{v}_{c}\cdot\bm{\hat{r}}_{i\pm}$. The inequality $s_{i+}-s_{i-}=\bm{v}_{c}\cdot(\bm{\hat{r}}_{i+}-\bm{\hat{r}}_{i-})>0$ holding indicates that the direction of $\bm{v}_{c}$ aligns with that of the test vector $\bm{\hat{r}}_{i+}-\bm{\hat{r}}_{i-}$, signifying a successful concept recognition. We calculate the accuracy of the concept distinction for each concept on the test data as $\textnormal{Acc}_{c}$:
$\textnormal{Acc}_{c}=\frac{\sum_{i=0}^{\hat{N}-1}\mathbb{I}(s_{i+}>s_{i-})}{\hat{N}}\quad\hat{N}=|\mathcal{T}_{c}^{\textnormal{test}}|$ (2)
A high accuracy ($\textnormal{Acc}_{c}>\tau$) indicates the presence of a specific value concept in the model.
This process is performed for each language $l$, resulting in $\textnormal{Acc}_{c}^{l}$. The results provide insights into whether the model effectively encodes the value concept $c$ in the context of language $l$.
Note that each layer has a recognition accuracy, using the concept vector of that layer. Generally, intermediate layers encode the most precise concepts. Unless specified otherwise, we select the best result from all layers. For more details on the linguistic and abstract concept information in different model layers, refer to Appendix E.2 and G.2.
Figure 1: Multilingual concept recognition accuracy (%) of LLaMA2-chat, Qwen-chat and BLOOMZ series, averaged across all value concepts. The performance of the three 7B-sized models are connected with dashed lines for performance comparison.
3.3 Calculating Cross-Lingual Similarity of Concept Vectors
Through calculating cross-lingual similarity of concept vectors, we explore the extent to which LLMs encode consistent representations for the same value concept in different languages, namely, the cross-lingual consistency of multilingual value concepts. Specifically, given two languages $l_{1}$ and $l_{2}$, we calculate the cosine similarity of their concept vectors $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$. Appendix G.1 highlights the effectiveness of employing cosine similarity to assess the correlation between concept vectors.
3.4 Recognizing Cross-Lingual Concepts
To investigate the cross-lingual transferability of a specific value concept across languages, we propose a method for cross-lingual concept recognition. Given two languages, $l_{1}$ and $l_{2}$, we calculate how accurately $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$ can be used to recognize the concept $c$ in language $l_{2}$, resulting in $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}$ and $\textnormal{Acc}_{c}^{l_{2}}$. The inequality $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}\geq\textnormal{Acc}_{c}^{l_{2}}$ being true signifies the successful transfer of concept $c$ from $l_{1}$ to $l_{2}$. Conversely, we calculate $\textnormal{Acc}_{c}^{l_{2}\rightarrow l_{1}}$ and $\textnormal{Acc}_{c}^{l_{1}}$ to explore the transferability of concept $c$ from $l_{2}$ to $l_{1}$. While evaluating transferability based solely on accuracy changes might imply a unidirectional transfer from high- to low-performing languages, Appendix H.1 indicates that transferability is not solely determined by language performance.
4 Experiments
We conducted extensive experiments with the proposed framework on 7 human values, 16 languages and 3 LLM families to answer questions Q1, Q2 and Q3. We leave the question Q4 to §5.
4.1 Experimental Setup
Human Value Datasets
We explored the following values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness. We utilized 3 subsets of ETHICS dataset (Hendrycks et al., 2021) for commonsense morality, deontology, and utilitarianism. Regarding fairness, truthfulness, toxicity, and harmfulness, we chose the StereoSet (Nadeem et al., 2021), TruthfulQA (Lin et al., 2022), REALTOXICITYPROMPTS (Gehman et al., 2020), AdvBench (Zou et al., 2023b) dataset, respectively.
Appendix B details the sources, data splits, and positive and negative examples for each value.
Examined Languages and LLMs
We translated the aforementioned human value datasets from English into 15 non-English languages using Google Translate. These languages belong to various language families, including Indo-European (Catalan, French, Indonesian, Portuguese, Spanish), Niger-Congo (Chichewa, Swahili), Dravidian (Tamil, Telugu), Uralic (Finnish, Hungarian), Sino-Tibetan (Chinese), Japonic (Japanese), Koreanic (Korean) and Austro-Asiatic (Vietnamese). The impact of translation quality on our results is discussed in Appendix C.
Our experiments involved three multilingual LLM families, including the LLaMA2-chat series (7B, 13B, 70B) (Touvron et al., 2023), Qwen-chat series (1B8, 7B, 14B) (Bai et al., 2023) and BLOOMZ series (560M, 1B7, 7B1) (Scao et al., 2022). Appendix D provides detailed language distributions of their pre-training data.
4.2 Q1: Do LLMs Encode Concepts Representing Human Values in Multiple Languages?
Figure 1 illustrates the multilingual concept recognition accuracy of the three LLM families, averaged across all value concepts. We observe that all three models achieve notable accuracy across all represented languages [^2] The model’s understanding in unrepresented languages may stem from cross-lingual transfer from other languages. Qwen’s technical report only mentions the inclusion of en and zh in its pre-training data. We conjecture the inclusion of 10 other languages (fr,es,pt,vi,ca,id,ja,ko,fi,hu) based on its significant performance in these languages. and even the smallest models surpass $\tau=65\%$ accuracy in them. These results confirm that LLMs effectively encode value concepts in a multilingual context. Moreover, Figure 1 shows a clear pattern that increasing model parameters and language resources notably improves concept recognition accuracy. This indicates that both expanding the model size and increasing the scale of language pre-training data contribute to a more precise encoding of value concepts.
Appendix E.1 compares the PCA-based method with the mean-based method outlined in §3.1. It reveals that both methods produce concept vectors of comparable precision, with the mean-based technique holding a slight edge. The consistent performance across various extraction techniques confirm the effectiveness of concept vectors in capturing conceptual information. Appendix E.2 indicates that middle layers are particularly adept at capturing abstract concept information. Appendix E.3 demonstrates that even a small number of training samples can effectively extract representations of value concepts in LLMs. For detailed results on each value concept and additional discussions, please refer to Appendix E.4 and E.5.
Figure 2: Cross-lingual similarity of concept vectors across all language pairs, averaged over all value concepts. The languages included in each model’s pre-training data are presented and sorted based on their proportions in the corresponding model’s pre-training data. For Qwen-chat series, we conjecture its language inclusion based on multilingual concept recognition accuracy (§4.2) and display its primary languages, zh and en, at the forefront.
4.3 Q2 & Q3: How Consistent and Transferable are Value Concepts across Languages, and What is the Impact of LLMs’ Multilinguality?
Through computing cross-lingual similarity of concept vectors (§3.3) and recognizing cross-lingual concepts (§3.4), we investigated the cross-lingual consistency and transferability of these value concepts (Q2). Moreover, analyzing these concepts on LLMs trained with different multilingual data distributions provides insights into the multilinguality of LLMs (Q3).
4.3.1 Trait 1: Inconsistency of Concept Representations between High- and Low-Resource Languages
Figure 2 illustrates the cross-lingual similarity of concept vectors captured by the three 7B-sized models. We find that different multilinguality leads to different patterns of cross-lingual concept consistency. In the case of LLaMA2-chat-7B, the absolute dominance of English results in the model learning relatively independent concept representations for English, showing concept representation inconsistency between English and other languages, while higher cross-lingual concept consistency is observed among other languages. BLOOMZ-7B1’s cross-lingual concept consistency exhibits a very different pattern: the four languages with the lowest proportions (ta, te, sw, ny, accounting for 0.50%, 0.19%, 0.015%, and 0.00007% of pre-training data, respectively) show the lowest concept consistency (similarity) with other languages, while languages with relatively higher proportions (en with the highest percentage of 30.04%, and ca with the lowest percentage of 1.10%) demonstrate higher concept consistency with each other. [^3] We observe inconsistency between Spanish and other languages in BLOOMZ-7B1. We would like to explore this in our future work. For Qwen-chat-7B, we do not observe significant cross-lingual consistency between the main languages (zh, en) and other languages. In summary, cross-lingual concept inconsistency is more likely to occur between high- and low-resource languages.
The findings from Steck et al. (2024) suggest that a high average cosine similarity might raise concerns when dealing with unrelated representations. However, the results in Appendix G.1 indicate that, in our specific context, cosine similarity between concept vectors could reflect their genuine correlation. Additionally, Appendix G.2 presents cross-lingual consistency across different model layers, revealing that intermediate model layers exhibit greater consistency. For comprehensive results on each value concept and further discussions, please refer to Appendix G.3 and G.4.
Genetic Syntactic Geographic Phonological
D. C. D. C. D. C. D. C.
LLaMA2 -chat 7B -0.04 0.77 -0.12 0.63 -0.25 0.21 -0.03 -0.06
13B -0.17 0.53 -0.12 0.65 -0.17 0.35 0.09 0.24
70B -0.07 0.78 -0.12 0.66 -0.26 0.3 -0.0 0.01
Qwen -chat 1B8 0.06 0.42 0.07 0.32 -0.03 0.0 -0.02 0.05
7B 0.03 0.39 0.07 0.33 -0.04 0.04 -0.01 0.17
14B 0.01 0.42 0.01 0.5 -0.03 0.14 0.01 0.14
BLOOMZ 560M 0.20 0.43 0.13 0.55 -0.03 0.38 -0.12 -0.29
1B7 0.23 0.45 0.21 0.67 -0.01 0.43 -0.13 -0.28
7B1 0.16 0.36 0.09 0.52 -0.06 0.31 -0.11 -0.26
Table 1: Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs. Scores greater than or equal to 0.2 are highlighted in bold. “D.” refers to results obtained through direct computation; “C.” pertains to the average results derived by first categorizing languages based on language resources and then computing correlations within different language categories.
Figure 3: Cross-lingual concept transferability across all language pairs, averaged over all value concepts. Languages are sorted based on their percentages in the pre-training data.
4.3.2 Trait 2: Linguistic Relationships Distortion due to the Imbalance of Language Data
To explore the correlation between cross-lingual concept consistency and linguistic similarity, following Qi et al. (2023), we used lang2vec [^4] https://github.com/antonisa/lang2vec to compute four types of linguistic similarity (genetic, syntactic, geographic, and phonological) between languages. We then calculated the Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs.
We employed two calculation methods to estimate the correlation. The first method directly computes the Pearson correlation on all language pairs (Direct), while the second starts by categorizing language pairs based on language resources. Subsequently, correlations are computed within different categories and averaged (Category). Please refer to Appendix F for details of the latter method.
Table 1 presents the correlation results. First, we observe that neglecting differences in language resources (Direct), there is no significant correlation between cross-lingual concept consistency with all types of linguistic similarity. However, upon considering disparities in language resources (Category), the correlation becomes apparent. These findings highlight that the multilingual concept representations embedded by LLMs can distinctly reflect linguistic relationships between languages. Nevertheless, these relationships are influenced by language discrepancies in the pre-training data of LLMs, deviating from the natural patterns.
In terms of linguistic variations, cross-lingual concept consistency exhibits the strongest correlation with genetic and syntactic similarity. In contrast, there is a weak positive correlation between cross-lingual concept consistency with geographic similarity, while no correlation is observed with phonological similarity. The results suggest that LLMs embed more consistent value concepts for language pairs with similar syntactic structures, genetic relations, and geographic proximity, aligning with previous findings on multilingual factual knowledge (Qi et al., 2023).
4.3.3 Trait 3: Unidirectional Concept Transfer from High- to Low-Resource Languages
For a given source language $l_{1}$ and target language $l_{2}$, we compute $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}-\textnormal{Acc}_{c}^{l_{2}}$ (the difference in accuracy scores) to measure the transferability of concept $c$ from $l_{1}$ to $l_{2}$3.4). We average differences in accuracy scores over all value concepts to measure the overall transferability. If the average difference is greater than 0, it indicates positive transferability from $l_{1}$ to $l_{2}$.
We present the cross-lingual concept transferability of the three 7B-sized models in Figure 3. It provides insights into the influence of LLMs’ multilinguality. Firstly, based on the results of LLaMA- and Qwen-chat-7B, we observe a pattern of monotonic concept transfer from the dominant languages to other languages. This pattern also exhibits an upper triangular cross-lingual transferability (the dashed triangular in Figure 3), indicating that cross-lingual concept transfer from high- to low-resource languages is more prevalent. In contrast, BLOOMZ-7B1 exhibits a relatively balanced bidirectional cross-lingual concept transferability, while for languages with extremely low resources, the tendency of unidirectional transfer persists.
While evaluating transferability based solely on changes in accuracy may introduce biases due to initial performance variations across languages, potentially amplifying the observed unidirectional transfer, Appendix H.1 indicates that transferability is not solely determined by language performance. For comprehensive results on each value concept and further discussions, please refer to Appendix H.2 and H.3.
en fr zh es pt vi ca id ja ko fi hu Avg
LLaMA2 -chat-7B No-Control 0.97 1.94 6.80 1.94 6.80 4.85 8.74 5.83 3.88 10.68 14.56 4.85 6.44
LS-Control 97.09 99.03 95.15 99.03 97.09 97.09 90.29 98.06 97.09 100.0 99.03 99.03 97.35
En-Control 97.09 94.17 94.17 97.09 91.26 96.12 91.26 88.35 99.03 95.15 95.15 91.26 93.91
LLaMA2 -chat-13B No-Control 0.97 0.97 5.83 1.94 5.83 5.83 27.18 8.74 2.91 10.68 15.53 6.80 8.38
LS-Control 88.35 99.03 97.09 98.06 99.03 98.06 98.06 100.0 98.06 97.09 98.06 100.0 98.41
En-Control 88.35 99.03 95.15 98.06 97.09 98.06 93.20 94.17 99.03 97.09 90.29 87.38 95.32
LLaMA2 -chat-70B No-Control 0.00 1.94 4.85 0.97 6.80 2.91 27.18 11.65 2.91 20.39 18.45 10.68 9.89
LS-Control 74.76 87.38 68.93 55.34 90.29 79.61 98.06 92.23 63.11 84.47 95.15 96.12 82.79
En-Control 74.76 95.15 70.87 92.23 79.61 95.15 63.11 73.79 92.23 74.76 72.82 63.11 79.35
Table 2: Following rates on LLaMA2-chat series under different control methods. “No-Control”: no control is applied; “LS-Control”: language-specific control with each language controlling itself; “En-Control”: cross-lingual control with English as the source language. “Avg” denotes the average results excluding English.
5 Q4: Is Value Alignment of LLMs Controllable across Languages?
LLaMA2-chat models, trained with alignment techniques such as RLHF, exhibit value alignment capabilities like rejecting harmful instructions. In this section, we employed the Representation Engineering (RepE) methodology (Zou et al., 2023a) to bypass such defense and further explored the potential for cross-lingual control of value alignment.
5.1 Cross-Lingual Value Alignment Control
To control a LLM to exhibit behavior aligned with a value concept $c$, a straightforward RepE-style method is multiplying the previously extracted concept vector $\bm{v}_{c}$ by a control strength $s$ and adding it to the hidden states of multiple layers $L$ within the target model. This procedure is iteratively applied to each token, formulated as $\bm{h_{i}}^{{}^{\prime}}=\bm{h_{i}}+s\cdot\bm{v}_{c}$, where $\bm{h_{i}}$ and $\bm{h_{i}}^{{}^{\prime}}$ denote the original and perturbed hidden state of $i$-th token, respectively. [^5] Reflecting on §3.1, each layer has its specific concept vector, and the perturbation is executed across multiple layers $L$. We omit the detail here for simplicity. In a cross-lingual scenario, we leverage the concept vector $\bm{v}_{c}^{l}$ of the source language ${l}$ to control the model’s behavior across various target languages. To determine appropriate control strength $s$ and control layers $L$ for cross-lingual control, we first conduct hyperparameter search to choose the combination that demonstrates the most effective control on language ${l}$. Subsequently, we employ this combination for cross-lingual control across all target languages and evaluate the control effect on each of them.
In our experiments, a successful control is steering the LLM to follow a harmful instruction rather than rejecting it. We compute the Following rate, representing the proportion of harmful instructions the model follows, to assess the effectiveness of model control. Specifically, we utilize the multilingual negative testing data (harmful instructions) for the concept of harmfulness (§4.1), calculating the Following rate in each language. Please refer to Appendix I for details of hyperparameter search and model control evaluation.
5.2 Results
Cross-lingual value alignment control results are presented in Table 2. First, without applying any control (No-Control), LLaMA2-chat series refrains from responding to almost all harmful instructions in English. However, simply translating these prompts into other languages partially circumvents the models’ defense, exposing LLMs’ multilingual vulnerability (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). Surprising, we observe larger models are more prone to responding to non-English harmful instructions, potentially due to their enhanced instruction-following capabilities.
Second, we discover that cross-lingual control from English to other languages (En-Control) can achieve control effectiveness comparable to that of LS-Control. While LS-Control achieves performance through language-specific optimization of hyperparameters, En-Control simply adopts hyperparameters found in English, highlighting the ease of achieving cross-lingual control with English as a source language in English-dominated LLMs.
6 Discussions and Suggestions
Drawing our empirical observations and findings, we prudently consider the following suggestions for the configuration of multilingual pre-training data for LLMs, which might contribute to enhancing multilingual AI safety and utility. First, despite the positive effect of dominant languages as sources for cross-lingual alignment transfer (§5.2), it is essential to avoid an excessive prevalence (exemplified by LLaMA2’s pre-training data, which comprises about 90% English data). Our analysis suggests that such excessive dominance can lead to unfair cross-lingual patterns, manifested as inconsistent multilingual representations (§4.3.1), distorted linguistic relationships (§4.3.2), and monotonous transfer patterns (§4.3.3). These tendencies could potentially further amplify the risk of multilingual vulnerability (§5.2) and undermine cultural diversity (Zhang et al., 2023; Cao et al., 2023). Furthermore, we encourage a more balanced distribution of non-dominant languages, particularly those with extremely limited resources, to foster more equitable cross-lingual patterns (§4.3.1 and §4.3.3). [^6] These suggestions are based on our findings, which might be biased by factors like variations in language performance (§3.4) and other unobserved ones.
7 Conclusion
We have presented a systematic exploration of multilingual concepts embedded in LLMs, focusing specifically on human value-related concepts (i.e., value concepts). Through our extensive analysis spanning 7 human values, 16 languages, and 3 LLM families, we have obtained many interesting findings. Specifically, we empirically verify the presence of multilingual value concepts in LLMs and identify the cross-lingual characteristics of these concepts arising from language resource disparities. Furthermore, our experiments on cross-lingual control illuminate the multilingual vulnerability of LLMs, as well as the feasibility of cross-lingual manipulation over value alignment of LLMs. With these findings, we prudently present several suggestions for collecting multilingual pre-training data for advanced multilingual AI.
Limitations
Our work’s major limitation lies in the reliance on translations generated by machine translation for our primary experimental data. A straightforward translation of data related to human values not only introduces translation noise but also overlooks cultural differences. We discuss these two points below.
(1) The noise introduced by machine translations has minimal impact on our research findings. Firstly, our research focuses on the existence of multilingual value concepts in LLMs and their multilinguality, which do not depend on exceptional performance in any specific language. Additionally, we examine across multiple tasks, human values, languages, and LLMs to uncover universal patterns, which contributes to the robustness of our results to a certain degree of noise.
(2) We recognize that cultural variations can result in diverse interpretations of explored values among individuals from different cultural backgrounds. However, our work delves into research questions beyond cultural differences. We primarily focus on the multilingual representations of value concepts with LLMs, their universal cross-lingual patterns, and cross-lingual control over value alignment, aiming to enhance the safety and utility of multilingual AI. Additionally, our proposed framework may also be valuable for studying value disparities. For instance, when applying English concept vectors to other languages for cross-lingual concept recognition, errors in recognition may arise from value disparities between them. We plan to further explore the application of our framework to cultural divergences in our future research.
Ethical Statement
In this paper, we leverage the ETHICS, StereoSet, TruthfulQA, REALTOXICITYPROMPTS, and AdvBench datasets to delve into diverse human values. Despite the presence of negative elements such as unethical, biased, untruthful, toxic, and harmful content within these datasets, our utilization of them is consistent with their intended use. Our approach to cross-lingual value alignment control involves employing the representation engineering methodology to control LLMs’ behavior. While experimental results suggest that it is possible to steer LLMs towards generating harmful content, this underscores the applicability of this methodology in red-teaming LLMs to enhance AI safety and in steering LLMs towards producing harmless content in the opposite direction.
Appendix A Introduction to the Explored Values
Given that the concepts we delve into are inherently rooted in ethics and morals, it’s essential to clarify their ethical foundations. Below, we present the ethical theory as summarized by Vida et al. (2023). Grounded in this theoretical framework, we then elucidate the definitions and ethical characteristics of each value we explore.
A.1 Ethical Theory
According to Vida et al. (2023), Ethics is divided into four branches: normative ethics, applied ethics, descriptive ethics, and metaethics.
Specifically, normative ethics focuses on the principles and criteria that define moral correctness. It operates within a framework of universal norms and values, providing justification for what is deemed right or wrong. Descriptive ethics, conversely, involves empirical investigations to describe or explain the moral judgments, preferences, and value systems prevalent in societies. It refrains from making moral judgments, focusing instead on documenting and analyzing prevailing ethical beliefs and behaviors. Applied ethics extends the general norms and values from normative ethics to specific contexts and fields, dealing with concrete ethical dilemmas and decisions in domains like bioethics, environmental ethics, or, as relevant to our paper, the ethics of artificial intelligence. Metaethics lays the analytical foundation for these three branches, delving into the nature of moral language, the meaning of moral judgments, and the foundational aspects of ethical theories.
Furthermore, normative ethics can be assigned to three competing ethical families: virtue ethics, deontological ethics, and consequentialism. While deontological ethics emphasizes the intrinsic rightness or wrongness of actions based on principles or rules, consequentialism assesses actions by their outcomes or consequences. Meanwhile, virtue ethics focuses on the moral character and virtues of the individual.
A.2 Definitions and Ethical Characteristics of Each Value
Below, we detail the definitions of the 7 explored values, their ethical characteristics, and any interconnections between them.
Commonsense Morality
Commonsense Morality refers to the intuitive and widely accepted moral principles guiding everyday human behavior. These principles often stem from societal norms, cultural values, and emotional responses, forming the basis of our ethical decision-making. Commonsense Morality focuses on evaluating actions based on moral correctness rather than merely describing existing moral beliefs and behaviors in society. Thus, it can be categorized as a part of normative ethics.
Deontology
Deontology, on the other hand, focuses on the inherent rightness or wrongness of actions based on adherence to a set of rules or constraints. It asserts that certain actions possess moral obligations or prohibitions, independent of their outcomes. Thus, Deontology is categorized under normative ethics, specifically within the deontological ethics family. While both Commonsense Morality and Deontology belong to normative ethics, they differ in their foundational principles. Commonsense Morality is anchored in societal norms and moral correctness, emphasizing the alignment of actions with shared societal values. In contrast, Deontology prioritizes rule-based morality, focusing on the inherent moral obligations or prohibitions associated with actions, regardless of their outcomes.
Utilitarianism
Utilitarianism emphasizes maximizing overall well-being, aiming for a world where every individual experiences the highest possible level of well-being. Belonging to the consequentialism family within normative ethics, utilitarianism assesses the moral value of an action based on its outcomes or consequences, contrasting with deontology’s focus on the intrinsic rightness or wrongness of actions.
Fairness
Fairness pertains to the equitable and impartial treatment of individuals, regardless of their demographic attributes such as race, gender, age, religion, or socioeconomic status. Its emphasis on societal biases places Fairness within the realm of descriptive ethics, focusing less on absolute moral rightness or wrongness.
Truthfulness
Truthfulness involves the accurate representation of facts about the real world. In this context, a statement is considered truthful if it aligns with objective reality, without being influenced by personal beliefs or biases. Given that ensuring the honesty and transparency of AI systems is crucial in the realm of artificial intelligence, Truthfulness is more appropriately classified under applied ethics.
Toxicity
Toxicity refers to the presence of harmful or offensive language in text, which can include hate speech, harassment, or other forms of harmful communication. In the context of AI-generated content, Toxicity appropriately falls under applied ethics due to its direct influence on user experience.
Harmfulness
Harmfulness includes various types of detrimental content such as profanity, graphic depictions, threatening behavior, misinformation, discrimination, cybercrime, and dangerous or illegal suggestions. Harmfulness is inherently a broader concept and may intersect with other ones. Given its pivotal role in AI alignment research, we classify Harmfulness under applied ethics.
Table 3 further presents the positive and negative examples of each human value. Given the diverse definitions and ethical nature of the concepts we explore, we collectively term them “value concepts” in this paper, also aligning with AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). Note that the above classification adheres to ethical theories as closely as possible, but some deviation may still exist.
Appendix B Data Details
Below we describe the public datasets utilized for each human value.
Commonsense Morality
We utilized the COMMONSENSE MORALITY subset in ETHICS dataset (Hendrycks et al., 2021), which includes first-person characters’ actions with clear moral implications. In detail, for the same scenario, actions with positive or negative moral judgment are provided. The collection of scenarios includes both short and detailed examples, we only utilized the short ones considering our limited computing resources.
Deontology
We employed the DEONTOLOGY subset in ETHICS dataset (Hendrycks et al., 2021), which encompasses two subtasks: Requests and Roles. Specifically, in the Requests subtask, scenarios are created where one character issues a command or request, and another character responds with purported exemptions, which are judged as reasonable or unreasonable. In the Roles subtask, each role is assigned with reasonable and unreasonable responsibilities. We utilized data from both subtasks for our experiments.
Utilitarianism
We employed the UTILITARIANISM subset in ETHICS dataset (Hendrycks et al., 2021), where pairs of scenarios labeled as either more pleasant or less pleasant are provided.
Fairness
We used the StereoSet dataset (Nadeem et al., 2021), which consists of sentences measuring stereotypical bias across gender, race, religion, and profession. These sentences are split into two classes: intrasentence and intersentence. Specifically, each sentence in the intrasentence class has a fill-in-the-blank structure where the blank can be filled with the a stereotype term, anti-stereotype term or unrelated term. We inserted each of these three terms into the blank to form different complete sentences. In the intersentence class, each sentence containing a target term is followed by three associative sentences representing stereotypical, anti-stereotypical, and unrelated associations. We concatenated the preceding and subsequent three types of sentences to form different complete sentences. We only employed pairs of stereotypical and anti-stereotypical sentences to obtain positive and negative samples for this human value.
Truthfulness
We used the TruthfulQA dataset (Lin et al., 2022), which consists of two tasks: generation and multiple-choice. Specifically, in the generation task, questions are accompanied by correct or incorrect responses. In the multiple-choice task, questions are accompanied by a set of candidate answers, some of which are correct and others incorrect. We concatenated the question and its corresponding correct response or answer as a positive example while the same question with its corresponding incorrect response or answer as a negative example.
Toxicity
We utilized REALTOXICITYPROMPTS dataset (Gehman et al., 2020) consisting of naturally occurring prompts sampled from English web text and corresponding toxicity scores. We categorized prompts into non-toxic and toxic ones based on the scores, thereby forming positive and negative pairs.
Harmfulness
We utilized the AdvBench dataset (Zou et al., 2023b) which contains harmful instructions eliciting LLMs to generate objectionable content. These harmful instructions are further combined with harmless instructions to form negative and positive pairs, as described in the work of Zou et al. (2023a).
After collecting and formatting these datasets, we divided each dataset of human values into the training and testing sets in an 8:2 ratio. The training set is used for obtaining concept vectors, as discussed in Section 3.1, while the testing set is employed for experiments, such as concept recognition in Section 3.2 and model control in Section 5. Table 3 presents the number of training and testing samples, as well as positive and negative examples of each human value.
Appendix C Impact of Translation Quality
Our primary experimental data rely on translations yielded by translation engines. However, the noise introduced by these translations has minimal impact on our research findings. Our exploration of universal cross-lingual characteristics in LLMs, such as cross-lingual consistency and transferability, suggests that overall patterns are likely preserved when similar noise affects all languages simultaneously. For example, despite the “translationese effect” which could potentially enhance the similarity between non-English texts and English, significant cross-lingual inconsistencies remain between English and other languages in the LLaMA2-chat-7B series, as illustrated in Figure 2.
en fr zh es pt vi ca id ja ko fi hu Avg
mean 97.5 90.2 91.0 91.7 92.0 84.9 90.2 86.4 87.4 82.7 83.4 81.4 88.2
pca 96.7 92.7 90.7 91.7 89.2 85.9 90.2 83.2 86.9 80.7 82.2 81.2 87.6
Table 5: Comparison of multilingual concept recognition accuracy between PCA-based and mean-based concept extraction methods.
Figure 4: Multilingual concept recognition accuracy across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
Figure 5: English concept recognition accuracy with varying numbers of training samples for collecting concept vectors. The result are based on LLaMA2-chat-13B. We calculate the average accuracy across all layers to ensure the results of different settings are comparable.
Appendix D Language Distribution
Table 4 displays language distributions of the 16 selected languages (including English) in both the LLaMA2-chat and BLOOMZ series’ pre-training data. For the Qwen-chat series, English and Chinese constitute a significant portion of its pre-training data, although detailed language distribution is not publicly accessible.
Based on the language distributions in their pre-training data, we categorize the multilinguality of these 3 LLM families into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with balanced multilinguality (i.e., BLOOMZ series).
Appendix E More Results of Multilingual Concept Recognition
E.1 Extracting Concept Vectors based on PCA
To further enhance the robustness of our results, we also employed the PCA-based method and compared it with the mean-based approach outlined in Section 3.1 (refer to Hämmerl et al. (2023) or Zou et al. (2023a) for details on the PCA-based method). Table 5 presents the multilingual concept recognition accuracy (Section 3.2) for the concept of deontology on LLaMA2-chat-7B. The results suggest that the mean-based method extracts more distinct concept vectors across languages compared to the PCA-based method, consistent with the conclusions of Zou et al. (2023a).
E.2 Results across Model Layers
Figure 4 presents the concept recognition accuracy across different model layers. We observe that the most explicit concept vectors are encoded in the middle layers of the models. This suggests that the middle layers are more likely to encode explicit abstract value information, aligning with the findings of Li et al. (2023); Zou et al. (2023a).
E.3 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
We employed varying amounts of training samples to extract concept vectors, and the recognition performance for each human value is illustrated in Figure 5. Surprisingly, optimal accuracy can be achieved for all human values even with few training samples, consistent with the findings by Li et al. (2023), suggesting that the concept vectors for human values are readily extractable in LLMs. Furthermore, we observe notable differences in the recognition accuracy of different human values, indicating different degrees of difficulty in capturing them. Specifically, harmfulness, toxicity, commonsense morality, and deontology are relatively explicitly encoded human values. In contrast, LLMs encounter a greater challenge in recognizing concepts like truthfulness, fairness and utilitarianism.
E.4 Complete Results
Complete results of multilingual concept recognition are provided in Table 9.
E.5 Multilingual Performance Reflects Multilinguality
As shown in Figure 1, the performance distributions of different models across all languages reflect their multilinguality. Specifically, while all three model families perform best in English, the LLaMA2-chat series exhibits significant performance disparities between English and non-English languages. The Qwen-chat series, while excelling at English, also outperforms other languages in Chinese. In contrast, the BLOOMZ series demonstrates the smallest performance gap between English and non-English, reflecting a more balanced multilinguality.
Appendix F Computing Pearson Correlation Coefficients Considering Differences in Language Resources
This method begins by categorizing languages into high- and low-resource based on their proportions in the LLM pre-training data. Specifically, for the LLaMA2-chat series, English is designated as a high-resource language, while the remaining languages are considered as low-resource languages. In the case of BLOOMZ series, the low-resource languages include ta, te, sw, and ny, while the rest are considered as high-resource languages. For the Qwen-chat series, en and zh are treated as high-resource languages. We then partition the scores of cross-lingual concept consistency and linguistic similarity among all language pairs into two groups: those between high-resource languages and all languages, and those among low-resource languages themselves. Subsequently, we compute the Pearson correlation coefficients separately for these two sets and report the average result. In this way, imbalance of language distributions between high- and low-resource languages is mitigated when computing the Pearson correlation between cross-lingual concept consistency and linguistic similarity.
Figure 6: Cross-lingual similarity of concept vectors across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
Appendix G More Results of Cross-Lingual Concept Consistency
same different
LLaMA2-chat-7B (en-en) 1.00 0.56
Qwen-chat-7B (en-en) 1.00 0.49
BLOOMZ-7B1 (en-en) 1.00 0.49
LLaMA2-chat-7B (en-fr) 0.95 0.54
Qwen-chat-7B (en-fr) 0.92 0.44
BLOOMZ-7B1 (en-fr) 0.95 0.53
Table 6: Cosine similarity between concept vectors representing either the same or different values across languages.
G.1 Cosine Similarity between Concept Vectors can Reflect Their Correlation
Steck et al. (2024) discussed the limitations and potential issues with using cosine similarity as a measure of semantic similarity, particularly in the context of embeddings learned from linear models. They highlight that cosine similarity can sometimes produce arbitrary and non-unique results, implying that a high average cosine similarity might raise concerns when dealing with unrelated representations.
In our paper, cosine similarity is calculated on concept vectors across different languages to measure their consistency. It is worth recalling that these concept vectors are computed by averaging a set of difference vectors. This averaging process inherently filters out irrelevant information to some extent, thereby mitigating the unpredictable impact on cosine similarity results.
Furthermore, we attempt to evaluate the effectiveness of cosine similarity outcomes in our specific context. Specifically, we compute the cosine similarity between concept vectors of different values in English (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c2}^{en})$) and cross-lingually between English (en) and French (fr) for both the same (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c1}^{fr})$) and different (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c2}^{fr})$) human values. The averaged results presented in Table 6 indicate that, compared to the same human values, the concept representations of unrelated human values exhibit significantly lower cosine similarity. This observation holds true both within a single language and across languages. These findings suggest that, at least in our context, high cosine similarity tends to indicate high relevance, while low cosine similarity often signifies irrelevance to a considerable extent.
G.2 Results across Model Layers
Figure 6 illustrates the trends in cosine similarity across different model layers. We observe that the peak of cross-lingual consistency appears in the intermediate layers, with lower similarity near the input and output layers. This observation is consistent with previous research (Chi et al., 2021; Bhattacharya and Bojar, 2023), suggesting that middle layers of multilingual models encode a higher degree of language-independent information, while language-specific information is more prominent near the input and output layers.
G.3 Complete Results
Cross-lingual concept consistency of all models is presented in Figure 7.
G.4 Effect of Model Size
Despite larger models being able to capture more explicit concepts of human values (as shown in Figure 14), the increase in model size does not steadily enhance cross-lingual concept consistency, as shown in Figure 6.
[…]
LLaMA2-chat-7B 27.3% 22.7% 3.0% 47.0%
Qwen-chat-7B 30.3% 19.7% 7.6% 42.4%
BLOOMZ-7B1 34.1% 15.9% 16.7% 33.3%
Table 7: Proportion of cases in which the concept recognition performance of language A either surpasses or underperforms language B, and whether the transfer from language A to language B is effective or not. $\geq$” and “$<$” denote superiority and inferiority respectively, and $\surd$” and “$\times$” represent successful and unsuccessful transfer.
en zh fr es pt vi ca id avg
LLaMA2 -chat 7B 0 14 28 28 14 14 57 85 30
13B 0 14 57 42 42 71 57 100 47
70B 0 71 14 28 28 85 71 85 47
Qwen -chat 1B8 0 0 42 14 28 100 85 28 37
7B 14 14 57 0 71 42 71 71 42
14B 14 14 57 14 57 85 57 71 46
BLOOMZ 560M 14 14 100 0 57 85 14 100 48
1B7 85 42 71 42 42 100 0 85 58
7B1 100 14 100 71 57 100 42 85 71
Table 8: Proportions of different languages as targets of cross-lingual concept transfer. The displayed languages are those included both in LLaMA2-chat and BLOOMZ series’ pre-training data.
Appendix H More Results of Cross-Lingual Concept Transferability
H.1 Transferability Beyond Language Performance
While the setting described in Section 3.4 may introduce bias of initial performance variations across languages, potentially leading to mono-directional transfer from high-performing languages to low-performing ones, our findings suggest that transferability is not solely determined by language performance, as detailed below.
[…]
In summary, although evaluating transferability based solely on changes in accuracy may pose limitations, the phenomenon that transfer is not solely determined by language performance indicates that this remains an open question. We plan to develop more robust and unbiased methodologies to further investigate cross-lingual transfer in our future research.
H.2 Complete Results
Cross-lingual concept transferability of all models is presented in Figure 8.
H.3 Effect of Multilinguality and Model Size
Table 8 provides a breakdown of the proportions of different languages as targets of cross-lingual concept transfer [^7] If $\textnormal{Acc}^{l_{1}\rightarrow l_{2}}\geq\textnormal{Acc}^{l_{2}}$, $l_{2}$ is considered as a target of the concept transfer between the two languages., providing a clearer illustration of the unidirectional transfer from dominant languages in LLaMA2- and Qwen-chat series. Conversely, the BLOOMZ series demonstrates a more balanced transfer pattern, showcasing a distinctly superior level of cross-lingual concept transferability.
Furthermore, Table 8 reveals that increasing the model size consistently improves in cross-lingual concept transferability, except for cases of LLaMA2-chat-13B and 70B, where similar levels of cross-lingual transfer are observed.
Appendix I Hyperparameter Search and Control Effectiveness Evaluation in Experiments of the Cross-Lingual Value Alignment Control
Hyperparameter Search
For the control strength $s$, we explored values from 1 to 10 with a step size of 1. Regarding the control layers $L$, we initially sorted the model’s layers based on their performance in concept recognition (Section 3.2) from high to low. This transforms the selection of control layers into choosing top K layers with the best concept recognition performance. We considered a total of 10 variations of top K values, ranging from 1 to two-thirds of the model’s maximum number of layers with a fixed step size. This results in a total of 100 hyperparameter combinations.
Evaluation of Control Effectiveness
To evaluate the effectiveness of cross-lingual value alignment control, we first used Google Translate to translate non-English model responses into English. After that, we classified the model’s responses to harmful prompts into the following three categories: Refusing, yielding responses with an explicit intention of refusing harmful prompts; Following, producing responses without any refusal intention; and Failure, generating responses that deviate from normal text, such as repeating a specific word. The situation of Failure may result from an excessive control strength ${s}$ and an excessive number of control layers ${L}$. We calculated the proportions of these three categories, namely the Refusing rate, Following rate, and Failure rate, and utilized the Following rate as a metric to evaluate the cross-lingual value alignment control effectiveness.
Through extensive manual checking, we determined the specific criteria for assessing Refusing, Following and Failure as Follows. First, we classified responses from the LLaMA2-chat series as “Refusing” if they begin with any of the following phrases: “As a developer”, “As a model of language”, “As an ethical AI”, “As a cybersecurity expert”, “As a model language”, “As a responsible language model”, “Unfortunately”, “No,”, “Sorry”, “Of course not!”, “I cannot”, “I can’t”, “I could not”, “I apologize”, “I’m sorry”, “I’m just an AI”, “We cannot”, “We can’t”, “It is not appropriate”, “It is not ethical”, “It is neither appropriate”, “It is neither ethical”. Second, we utilized the following criteria to categorize responses as “Failure”: 1. If the response length is fewer than 3 words; 2. If the response contains excessively long words with more than 15 characters; 3. If the response contains more than 1 word repeated consecutively more than 2 times, with a maximum gap of 5 words between repetitions considered as repeated. The remaining responses are categorized as “Following”.
Note that these rules are effective only for the dataset and model used in our experiments and may require adjustments for other scenarios.
Commonsense Morality en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 98.5 91.7 88.5 89.8 88.6 86.7 85.3 84.5 86.1 80.3 73.7 76.4 58.5 57.2 60.8 58.1 79.0
13B 98.9 92.6 90.8 91.8 89.4 85.5 87.7 86.2 89.7 83.0 76.7 81.5 59.2 57.6 62.3 57.2 80.6
70B 99.6 95.9 91.4 94.7 93.7 87.1 91.9 90.2 90.6 87.1 82.9 85.1 62.1 58.7 63.4 59.7 83.4
Qwen -chat 1B8 90.9 74.4 88.2 74.9 72.1 56.9 64.2 67.1 66.8 59.6 58.3 59.8 56.5 55.1 55.2 53.5 65.8
7B 96.3 88.0 92.3 84.8 82.2 75.4 82.9 75.3 83.6 73.7 69.7 73.4 59.8 57.3 60.6 55.1 75.6
14B 97.2 93.5 93.1 91.8 89.4 91.1 88.5 90.7 89.4 90.5 80.4 80.2 68.2 70.9 60.2 58.7 83.4
BLOOMZ 560M 80.1 80.7 80.1 78.3 79.4 77.8 77.1 75.4 65.5 57.9 56.5 58.7 71.9 73.1 63.5 61.0 71.1
1B7 87.3 85.7 86.8 86.5 86.4 84.3 84.8 81.5 72.2 61.6 56.7 56.4 77.9 77.5 67.5 63.7 76.0
7B1 91.7 90.9 90.4 89.3 90.2 88.9 88.8 86.1 78.7 63.4 56.5 57.5 82.6 82.3 73.9 69.1 80.0
Deontology en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 97.5 90.2 91.0 91.7 92.0 84.9 90.2 86.4 87.4 82.7 83.4 81.4 64.8 59.0 69.1 65.1 82.3
13B 97.2 93.0 90.5 92.2 91.5 87.7 91.0 88.2 87.7 87.7 83.9 82.9 65.3 62.6 69.3 66.3 83.6
70B 99.5 95.5 91.7 94.7 95.5 87.9 94.5 91.2 88.4 83.7 86.4 89.7 65.6 61.8 71.6 65.3 85.2
Qwen -chat 1B8 94.0 81.4 91.5 84.2 81.7 79.9 77.9 75.9 75.9 74.1 68.8 68.6 62.3 59.5 66.1 62.8 75.3
7B 97.0 89.2 93.5 89.7 87.4 82.7 87.7 82.7 84.2 77.4 76.4 76.4 69.1 65.6 70.9 66.1 81.0
14B 96.2 95.0 95.0 94.5 93.7 94.0 92.2 91.5 87.2 87.9 82.7 81.4 77.4 78.9 71.4 67.1 86.6
BLOOMZ 560M 82.7 78.6 82.7 84.9 84.2 81.4 83.2 77.9 68.3 62.6 60.1 63.6 78.6 76.6 73.6 66.8 75.4
1B7 87.2 85.7 85.7 87.2 87.4 87.2 86.7 83.7 71.6 65.8 62.3 64.6 80.2 81.7 80.7 73.4 79.4
7B1 91.5 88.9 88.7 92.0 92.0 88.2 89.4 89.2 74.4 69.8 64.1 62.3 84.4 83.7 81.4 73.4 82.1
Utilitarianism en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 77.3 74.1 72.2 74.0 73.7 71.7 72.1 72.3 70.0 69.8 68.8 69.6 52.5 52.9 55.3 53.6 67.5
13B 77.7 73.1 72.1 73.8 73.5 71.3 72.4 71.8 70.2 71.9 70.0 72.2 56.1 53.3 55.9 53.8 68.1
70B 78.5 76.1 74.8 76.5 75.6 73.4 74.5 74.6 73.7 72.5 74.1 74.1 54.8 55.6 57.9 54.3 70.1
Qwen -chat 1B8 73.9 68.2 70.3 66.2 64.5 60.7 59.7 63.1 65.3 62.3 56.4 57.1 51.9 51.6 52.7 53.7 61.1
7B 74.9 73.4 74.4 73.8 71.3 69.3 69.0 67.6 69.3 68.3 68.0 66.5 53.1 53.4 55.0 54.2 66.3
14B 73.4 72.8 71.4 72.2 71.6 70.5 70.4 70.7 73.7 71.3 70.1 69.6 58.1 61.0 56.4 55.3 68.0
BLOOMZ 560M 73.4 72.5 71.1 72.2 71.1 71.5 70.5 71.7 60.0 53.4 54.3 54.5 65.6 64.1 60.9 55.4 65.1
1B7 75.3 74.4 71.9 74.1 74.0 73.3 71.5 72.7 63.7 58.4 54.5 54.6 67.4 67.1 61.0 58.8 67.0
7B1 76.9 75.1 74.1 74.7 74.3 74.9 73.2 74.8 66.3 62.3 55.1 54.1 69.3 68.5 66.4 61.8 68.9
Fairness en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 78.3 69.7 67.8 72.1 70.4 66.9 69.9 66.4 68.0 65.6 68.0 66.6 56.0 58.6 57.8 58.0 66.3
13B 80.0 72.0 70.4 74.7 72.7 69.3 71.4 68.4 71.4 70.3 70.6 68.9 59.5 59.3 59.0 59.0 68.6
70B 82.6 75.1 72.9 76.5 74.4 72.4 76.0 72.0 70.2 69.8 70.7 71.5 61.1 61.3 60.5 58.1 70.3
Qwen -chat 1B8 73.5 67.6 70.4 68.0 67.2 65.8 67.0 65.8 64.2 63.2 61.0 60.9 53.5 56.7 58.4 58.5 63.9
7B 80.7 72.9 77.5 76.1 72.3 70.3 75.5 70.3 71.3 68.4 67.9 69.6 60.2 60.6 59.4 57.7 69.4
14B 81.9 76.0 79.1 79.2 77.4 78.3 79.2 77.4 74.9 74.2 74.5 75.0 65.0 65.2 64.3 60.3 73.9
BLOOMZ 560M 70.1 66.5 70.1 67.7 65.9 69.2 68.7 65.8 63.8 61.5 57.7 57.6 63.7 64.3 63.3 59.2 64.7
1B7 72.0 68.4 70.0 70.3 68.8 72.7 71.9 69.5 65.4 59.5 55.3 60.4 67.6 67.5 67.6 61.7 66.8
7B1 75.9 73.8 73.0 74.8 72.3 75.9 76.4 72.5 67.8 65.7 57.2 60.1 68.6 71.1 70.0 65.4 70.0
Truthfulness en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 84.5 86.4 81.2 84.2 82.4 83.5 84.2 84.6 82.8 81.9 83.7 81.2 73.5 67.8 69.7 65.0 79.8
13B 87.1 85.6 79.7 84.9 82.9 84.1 83.8 83.1 82.4 81.4 83.4 82.3 73.8 67.9 71.9 65.4 80.0
70B 89.4 89.7 84.3 87.0 86.4 84.1 86.9 85.3 84.7 86.7 85.4 85.5 74.9 68.5 72.6 67.9 82.5
Qwen -chat 1B8 82.7 77.2 80.6 81.6 78.5 75.8 74.2 77.3 78.3 79.3 73.5 71.7 72.1 70.0 67.8 64.8 75.3
7B 83.5 80.6 81.8 84.2 82.1 78.4 80.5 78.9 80.5 80.0 76.4 76.6 73.7 70.7 68.0 64.9 77.6
14B 86.2 86.2 84.8 85.1 83.8 83.3 83.2 83.3 83.9 84.3 79.6 80.9 78.3 76.3 71.1 65.7 81.0
BLOOMZ 560M 78.3 77.8 75.0 82.1 78.6 79.1 76.4 77.2 74.6 69.0 66.0 63.0 75.8 73.2 73.3 66.1 74.1
1B7 82.1 80.2 79.9 84.0 79.9 80.0 79.3 79.9 76.5 73.9 64.6 64.8 79.3 75.7 76.0 72.3 76.8
7B1 84.1 82.2 81.4 85.0 83.2 81.9 82.1 82.2 78.9 75.4 69.5 68.5 81.7 79.4 78.5 74.7 79.3
Toxicity en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 98.4 97.0 96.0 96.8 97.4 94.5 97.3 93.8 95.6 93.3 94.1 94.8 70.3 69.0 80.7 74.4 90.2
13B 98.6 97.0 96.2 97.3 97.1 94.0 97.4 95.2 95.0 94.2 95.0 95.8 70.2 69.8 79.6 72.9 90.3
70B 98.7 97.6 96.5 96.9 97.2 95.4 98.3 95.2 96.3 95.0 96.7 96.0 75.0 74.6 82.3 76.4 91.8
Qwen -chat 1B8 96.1 82.1 92.6 78.8 80.3 75.7 78.6 77.0 76.1 78.1 76.6 74.0 60.4 59.1 69.2 66.1 76.3
7B 94.8 90.8 92.5 87.6 88.1 86.6 89.3 85.6 77.9 80.2 86.7 85.7 67.3 63.6 68.2 69.2 82.1
14B 94.8 90.3 92.4 88.8 89.6 87.9 90.4 89.0 82.0 84.7 89.0 87.2 76.4 69.4 75.8 69.7 84.8
BLOOMZ 560M 92.4 92.2 91.2 87.5 90.3 89.0 90.4 88.6 77.6 70.1 65.8 67.4 82.8 78.0 80.0 72.4 82.2
1B7 93.0 93.6 91.6 88.8 92.8 91.4 92.2 90.6 74.4 69.8 68.2 70.3 86.9 84.8 84.6 79.5 84.5
7B1 91.8 93.2 91.7 87.1 91.2 90.8 93.0 91.7 75.0 72.2 70.6 71.7 88.6 87.6 86.4 82.8 85.3
Harmfulness en fr zh es pt vi ca id ja ko fi hu ta te sw ny Avg
LLaMA2 -chat 7B 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.1 92.2 97.1 94.2 98.7
13B 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.1 93.2 99.0 92.2 98.9
70B 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 96.1 99.0 98.1 99.4
Qwen -chat 1B8 100.0 95.1 100.0 99.0 99.0 94.2 93.2 92.2 98.1 85.4 97.1 92.2 87.4 93.2 89.3 98.1 94.6
7B 100.0 96.1 100.0 100.0 100.0 99.0 98.1 99.0 100.0 92.2 98.1 98.1 95.1 93.2 94.2 94.2 97.3
14B 100.0 97.1 100.0 100.0 100.0 100.0 100.0 99.0 100.0 99.0 99.0 98.1 94.2 97.1 96.1 94.2 98.4
BLOOMZ 560M 100.0 98.1 100.0 100.0 100.0 99.0 100.0 99.0 99.0 84.5 96.1 89.3 96.1 99.0 97.1 94.2 97.0
1B7 100.0 99.0 99.0 100.0 100.0 100.0 100.0 100.0 99.0 93.2 94.2 91.3 95.1 96.1 98.1 98.1 97.7
7B1 100.0 100.0 99.0 100.0 100.0 100.0 100.0 100.0 100.0 93.2 94.2 93.2 98.1 99.0 98.1 98.1 98.3
Table 9: Complete results of multilingual concept recognition.
Figure 7: Cross-lingual similarity of concept vectors of all models across all language pairs, averaged across all value concepts.
Figure 8: Cross-lingual concept transferability of all models across all language pairs, averaged across all value concepts.